OpenStudy (anonymous):

how do you solve 1 - cos^2 (20) - cos^2 (70)

OpenStudy (anonymous):

hello

OpenStudy (ash2326):

cos^2+sin^2x=1 Hi jay cos x= sin (90-x) cos ^2 x=1- sin^2x cos 20= sin 70 so 1--sin^2 70-cos^2 70 1-(sin ^2 70+cos ^2 70) 1-1 0

OpenStudy (anonymous):

hm...

OpenStudy (anonymous):

why do youadd sin and cos ?

OpenStudy (ash2326):

cos 20 is equal to sin 70 ,I substituted that in the equation. we have cos^2x+sin^2x=1 , this is an identity i used it cos^2 70+sin2^ 70=1 so 1-1=0

OpenStudy (anonymous):

but originally if cos20 changes to sin 70 then the next sign - doesn't have to be same ? instead of +?

OpenStudy (ash2326):

see we have 1-(cos^2 20)-cos^2 70 cos 20 = sin 20 cos^2 20= sin^2 70 1-(sin^2 70)-cos^2 70 1-1

OpenStudy (ash2326):

it's - only i took - out of the bracket -(sin2^ 70+cos^2 70) 1- (1)

OpenStudy (anonymous):

then doesn't -(sin2^ 70+cos^2 70) 1- (1) -> this change it to 1 - - ( 1) ?? sorry i am such a burden